OpenStudy (anonymous):

The circle below is centered at the point (4, -3), and has a radius of length 3. What is its equation?

OpenStudy (anonymous):

THE EQUATION OF THE CIRCLE IS (x-h)square+(y-k)square=(r)square Here (h,k)=(4,-3) and r=4 therefore (x-4)square+(y+3)square=4 square i.e.x square+16-8x+y square+9+6y=0 x square+ y square-8x+6y+9=0